### v1

# Do Code LLMs Do Static Analysis?

Abstract
This paper investigates code LLMs’ capability of static analysis during code intelligence tasks such as code summarization and generation. Code LLMs are now household names for their abilities to do some programming tasks that have heretofore required people. The process that people follow to do programming tasks has long been understood to require static analysis. For example, human programmers navigate the call graph of large programs to comprehend the different parts of those programs. Education in programming includes static analysis under the assumption that better static analysis skills beget better programming. Yet while popular culture is replete with anthropomorphic references such as LLM “reasoning”, in fact code LLMs could exhibit a wholly alien thought process to humans. This paper studies the specific question of static analysis by code LLMs. We use three different static analysis tasks (callgraph generation, AST generation, and dataflow generation) and three different code intelligence tasks (code generation, summarization, and translation) with two different open-source models (Gemini and GPT-4o) and closed-source models (CodeLlaMA and Jam) as our experiments. We found that LLMs show poor performance on static analysis tasks and that pretraining on the static analysis tasks does not generalize to better performance on the code intelligence tasks.
Keywords:
static analysis code intelligence tasks large language modelsLLM reasoning
1 Introduction
Popular discussion around large language models (LLMs) is replete with anthropomorphic references such as LLM “reasoning” and “creativity” (Tian et al., 2024; Zhao et al., 2024). Even dialogue in scientific literature uses biological terms such as “attention” and “hallucination” to refer to parts of LLM architecture or behavior. The implication is that LLMs think and act like people, with human intentions and human thought processes. One may conclude that this implication is clearly hyperbole – many authors argue that impressive LLM abilities are in fact merely pattern matching across big data input (Wang et al., 2023) and/or carefully curated finetuning (Chen and Mueller, 2024). Yet the processes that LLMs follow are controversial and there is not yet sufficient empirical evidence on specific LLM behaviors to draw strong conclusions.
One group of LLM behaviors in the spotlight is in software development tasks. Code LLMs are language models specializing in these tasks such as generating code samples from natural language prompts, code comment generation, and translation from one programming language to another. Many programmers use code LLMs every day as a normal part of their workflow. For example, Google estimated in November of 2024 that over 25% of its new code is written by AI and later reviewed by engineers (arstechnica, 2024). The rapid rise of LLM-generated code has led to anxiety about the safety and security of programs containing that code (Mohsin et al., 2024; Wang et al., 2024a), privacy and other legal issues (He et al., 2024), and even the future of software engineering as a career (Rasnayaka et al., 2024; Tona et al., 2024). Knowing how code LLMs do coding tasks could shed light on these concerns.
Meanwhile, a fundamental part of what people do during programming tasks is static analysis. Static analysis is the process of examining code without executing it (Ayewah et al., 2008). Studies of human programmers have for decades shown how people build a mental model of program behavior through static tracing of three relationships in particular: function invocations (i.e., function call relationships), data flow among variables, and the hierarchical structure of code elements, such as how expressions, statements, and blocks are organized (Littman et al., 1987). A rich body of literature describes these and similar static analysis tasks which underpin development of static analysis tools built into popular IDEs (Luo et al., 2021).
This paper asks the question: do code LLMs do static analysis? Humans do static analysis as a core part of software development, so if code LLMs follow a process similar to humans, then one would expect code LLMs to also do static analysis. We conduct experiments in which we 1) observe the performance of LLMs on software development tasks, 2) observe the performance of those same LLMs on static analysis tasks, 3) finetune open-source LLMs to do static analysis tasks, and 4) observe if the finetuned LLMs’ performance on development tasks changes. One would expect improved performance on development tasks if the code LLMs are trained to do static analysis better.
We found that code LLMs generally do a poor job at static analysis tasks but have strong performance at coding tasks. We also found that finetuning LLMs with static analysis tasks improved models’ ability to do simple static analysis tasks, but did not improve models’ ability on more complicated static analysis and coding tasks. We found these results over two open-source (CodeLlaMA, jam) and two closed-source models (GPT-4o, Gemini) and two programming languages (Java, C/C++). We studied three static analysis tasks (call graph, data flow graph, and abstract syntax tree generation) and three programming tasks (code summarization, code translation, and code generation). Our findings contribute to the ongoing academic debate about how LLMs function and have implications for task generalization, training procedures, and the design of future LLMs for software engineering tasks.
2 Background and Related Work
This section discusses studies of program comprehension and the role static analysis plays in that comprehension process, code large language models, and studies of capabilities of those models.
2.1 Static Analysis in Program Comprehension
This section discusses the studies of program comprehension and static analysis for program comprehension.
Empirical Studies Program comprehension has been a target of empirical studies for a decade. Several empirical studies have been conducted to understand how programmers comprehend code. Littman et al. (1987) found that programmers use data flow for comprehending code. Pennington (1987) also found that programmers combine data flow and control flow to comprehend source code in their study with 80 professional programmers. Shneiderman and Mayer (1979) have similar conclusion that syntactic information is as important as semantic information. The way that programmers represent the code mentally can be referred as mental models (Siegmund, 2016; O’brien, 2003; Storey, 2005; Harth and Dugerdil, 2017)Harth and Dugerdil (2017) classified this into three different periods: 1) classical period 2) optimistic period 3) pragmatic period. In the classical period, several different strategies on how programmers understand code were proposed. During the optimistic period, several tools and studies were developed and conducted. The pragmatic period focus more on measurement instead of creating new methods.
Static Analysis Various studies have established that three static analysis tasks in particular are crucial to program comprehension: 1) call graph analysis, 2) data flow analysis, and 3) Abstract Syntax Tree (AST) generation. Jiang et al. (2017) showed that programmers use call graphs to do change impact analysis. Konopka et al. (2018) conducted an empirical study with novice and intermediate programmers and showed the importance of using data flow analysis for program comprehension tasks. Another study conducted by  Maalej et al. (2014) also showed that some programmers drew the data flow and control flow chart during program comprehension. Hovemeyer et al. (2016) used ASTs to analyze students’ submission to help students on their programming tasks. Overall, these show that programmers rely on static analysis to do program comprehension tasks.
Static analysis helps programmers analyze code base without the need to execute the source code and any input. Different theories on static analysis have been proposed. Park et al. (2021) proposed the parametric static analysis to theorize the static analysis. Static analysis helps programmers in different aspects of software developments (Vassallo et al., 2020; Beller et al., 2016)Vassallo et al. (2018) found that developers use the warning message from automatic static analysis tools to find the bugs in the source code. Ayewah et al. (2008) combined the static analysis tools for bug detection. Static analysis can be used to aid the program comprehension (Eisenbarth et al., 2001). In the study conducted by Tymchuk et al. (2018), they showed that static analysis tools can also be served as an educational tool. Singh et al. (2017) showed that static analysis tools can reduce the workload of code review.
2.2 Code Large Language Models
Code LLMs have been the trend of in the recent year (Zhang et al., 2023; Jiang et al., 2024a; Chen et al., 2025) and have been applied to several different software engineering tasks (Nam et al., 2024; Gu, 2023; Zheng et al., 2025; He et al., 2025; Zhong and Wang, 2024). In this section, we discuss the history of code LLMs. Specifically, we focus on how LLMs are trained and how the source code is represented during different stages. Then, we discuss the application of LLMs in software engineering and studies on the capabilities of code LLMs.
History Sun et al. (2024a) divided the history of code LLMs into three groups: 1) neural language models for code 2) code pre-trained models 3) code LLMs. Natural language models represent the code as the natural language. Different methods have been proposed to represent code. Alon et al. (2019) encode each code token as a word vector. Zhang et al. (2019) showed that ASTs can better encode the information in the code. LeClair et al. (2020) applied graph neural network to represent code for code summarization. Code pre-trained models pretrain the language models with huge amount of source code and finetuned it with downstream tasks. Wang et al. (2021) pretrained the models with encoder-decoder transformer architectures. Li et al. (2023b) proposed a decoder architecture only model for pretrainig. More recently, the attention shift toward scaling up the models and in-context learning. The models include open source models Codellama (Roziere et al., 2023) and several closed source models (Brown et al., 2020; Anil et al., 2023).
Applications LLMs have been applied to different domains of software engineering tasks (Zheng et al., 2025, 2023; Jin et al., 2024)Sarsa et al. (2022) and Sun et al. (2023b) examine the capabilities of LLMs on code summarization. Khan and Uddin (2022) applied LLMs to generate the summarization. Koziolek et al. (2024) leveraged the LLMs and retrieval augmentation for code generations. Different evaluation methods have also been studied (Liu et al., 2024c, b)Yuan et al. (2024) proposed multi-agent systems for code translation. Pan et al. (2023) applied decoder-only LLMs to translate different programming languages into Python. In addition to program comprehension tasks, Xue et al. (2024) used LLMs to generate the test case for FinTech software testing. Lu et al. (2024) use LLMs for vulnerabilities detection with in-context learning. Yin et al. (2024) and a LLM-based framework for automatic program repair. Overall, LLMs are widely used in the software engineering.
Capability Studies Several studies on LLMs capability have been published (Yang et al., 2024; Li et al., 2024b; Jiang et al., 2024a; Fan et al., 2024)Yang et al. (2024) divided the capabilities of code LLMs into three different ares 1) programming skills 2) complex reasoning 3) structure knowledge. Li et al. (2022) showed that LLMs defeat around half of the programmers during programming competitions. Several studies have showed the promising results on using multiple LLMs as an agent for code generation (Liu et al., 2023a, b; Talebirad and Nadiri, 2023)Li et al. (2023a) showed that LLMs can do complex reasoning on debugging. Chen et al. (2022) showed that LLMs can reason the output of the code for several math problems. Although several papers showed promising results on reasoning, several studies have also showed that LLMs can only do reasoning when the code is simple (Liu et al., 2024a; Chen et al., 2024)Madaan et al. (2022) and Wang et al. (2022) showed the ability of LLMs for structure prediction after few-shot learning.
3 Experimental Design
Our experiments center around three topics: 1) we observe LLM performance on code development tasks, 2) we observe LLM performance on static analysis tasks in baseline, in-context learning setting, and finetuned setting, and 3) we finetune LLMs to do static analysis and determine if those models’ performance on code development tasks increases. This section covers our research questions, methods to answer those questions, and experimental settings.
We divide our experiments across open-source and closed-source LLMs. For open-source LLMs we can closely control experimental variables and conduct finetuning. For closed-source LLMs we can access the largest available models, but we give up control over many variables and the ability to finetune.
3.1 Research Questions
Our research objective is to determine if LLMs that do code development tasks also “know how” to do static analysis and to determine if models’ ability to do static analysis is connected to the models’ ability to do code development tasks. We ask the following Research Questions (RQs):
  RQ1
What is the performance of open-source LLMs on code development tasks when finetuned for those tasks under recommended conditions?
  RQ2
What is the performance of closed-source LLMs on static analysis tasks as is?
  RQ3
What is the performance of open-source LLMs on static analysis tasks when finetuned for those tasks?
  RQ4
What is the performance of open-source LLMs on code development tasks when finetuned for those tasks, after they have been finetuned for static analysis?
The rationale for RQ1 is to establish a baseline in a controlled environment of how well representative LLMs perform on code development tasks with common datasets and settings. Using open-source LLMs allows us to avoid data contamination issues that are present with these datasets in commercial LLMs (Jiang et al., 2024b). The rationale behind RQ2 is that state-of-the-art commercial LLMs are already known to perform code development tasks well (Zheng et al., 2025), but we do not yet know how well these models perform static analysis tasks. The rationale behind RQ3 is that LLMs may be able to perform static analysis tasks if they are explicitly trained to do so, and the availability of open source models provides us the opportunity to do this training. Finally, the rationale behind RQ4 is that models may do better at code development tasks if they can also do static analysis tasks because static analysis tasks are often a part of code development tasks.
3.2 Code Development Tasks
In this section, we discuss the three different code development tasks that we use i.e. code summarization, code translation, and code generation and the datasets for each task. We follow the instructions from the dataset because our idea is to show whether LLMs follow what human programmers do for code development tasks. We use C/C++ and Java as an example.
Code Summarization Code summarization is the task of writing natural language descriptions of a section of code (Haiduc et al., 2010). Code summarization is important because it is the core of automatic documentation generation and related tool support for helping programmers understand code; the history of research efforts in code summarization has been chronicled in recent surveys (Zhang et al., 2024b, 2022). We study this task in this paper because code summarization has become an advertised feature of code LLMs (Sun et al., 2024b) and because different studies show how people do static analysis in order to do code summarization (Rodeghero and McMillan, 2015; Stapleton et al., 2020). In this paper, for Java we use the “170k” dataset by Su and McMillan (2024). That dataset has its origins in a study of code summarization dataset recommendations by LeClair and McMillan (2019) and further vetted by (Bansal et al., 2021). A major advantage to the dataset is that it’s test set corresponds to a set held out of training in the jam model, which strongly avoids data contamination concerns (see Section 3.4). For C/C++ we use a dataset by Liu et al. (2021) which followed similar best practices (e.g., (LeClair and McMillan, 2019; Allamanis, 2019)).
Code Translation Code translation is the task of writing a program in one language that has the same functionality as a program in another programming language. Code translation can help program comprehension by in converting pseudo code into machine readable code (Holt and Turanski, 1960) and translate less-readable programming language into a more human-readable programming language (Hong and Ryu, 2024). Several studies have shown that code translation requires static analysis when done manually by people (Siddiqui et al., 2023; Feautrier, 1991). In this paper, we use code translation dataset in CodeXGLUE (Lu et al., 2021) as our dataset for Java. CodeXGLUE collects several datasets for code intelligence tasks. CodeXGLUE divides the datasets into code-code, code-text, text-code, and text-code. The code translation dataset that we use is collected from several public repositories with deduplication and clear train/test/validation separation. For C/C++, we use the C to Rust dataset in Huggingface website (Yu, 2023).
Code Generation Code generation is the task of generating source code that implements the functionality described in natural language. Code generation can be used as an educational tool to help novice programmers learn programming (Becker et al., 2023; Mosterman, 2006) and is widely used to improve the productivity of software developers (Ross et al., 2023; Xu et al., 2022). Static analysis is part of generating code in that an understanding of the static structure leads to higher code quality and few bugs (Barke et al., 2023; Nachtigall et al., 2019). In this paper, we used CONCODE datset (Iyer et al., 2018) from CodeXGLUE Lu et al. (2021) for Java. The dataset is collected from 33,000 repositories and they filtered out the methods without Javadoc as natural language description. Although the dataset has very clear train/validation/test split, we use separate half of the validation set as our testset because we do not find the ground-truth for testset at the time we download. For C/C++, we use Xlcost (Zhu et al., 2022). Xlcost is a dataset that contains seven different programming languages and natural language, though we only use the C/C++ component.
Table 1: Datasets and Models for Different Tasks in C/C++ and Java
Task Language Dataset Models
Name Source JAM CodeLlaMA
Development Code Summarization Java 170k  Su and McMillan (2024) x x
C/C++ Liu et al. (2021) Liu et al. (2021) x
Code Translation Java CodeXGLUE Lu et al. (2021) x x
C/C++ C to Rust  Yu (2023) x
Code Generation Java CONCODE Lu et al. (2021) x x
C/C++ Xlcost  Zhu et al. (2022) x
Static Analysis AST Generation Java funcom  LeClair and McMillan (2019) x x
C/C++  Haque et al. (2021)  Haque et al. (2021) x
Call Graph Generation Java funcom  LeClair and McMillan (2019) x x
C/C++  Haque et al. (2021)  Haque et al. (2021) x
Dataflow Graph Generation Java funcom  LeClair and McMillan (2019) x x
C/C++  Haque et al. (2021)  Haque et al. (2021) x
3.3 Static Analysis Tasks
In this section, we discuss three static analysis tasks in our experiments i.e. AST, data flow graph, and callgraph generation. We use different tools to generate the gold for each task. We use C/C++ and Java in our experiments.
Abstract Syntax Tree Generation An Abstract Syntax Tree (AST) is a tree structure that captures the syntactic information of the source code. In addition to being a key intermediate representation in code compilation, ASTs are used to localize software bugs (Neamtiu et al., 2005), comprehend source code (Sun et al., 2023a; Welty, 1997), and in software evolution (Neamtiu et al., 2005). Different studies have also shown that using ASTs in tool support and visualization can help program comprehension, such as when learning how to code (Agrahari and Chimalakonda, 2020; Egan and McDonald, 2014). In this paper, for Java we use dataset recommendation by LeClair and McMillan (2019). We use 52M to train jam (Su et al., 2023) and further extract 25k to train CodeLlaMA (Roziere et al., 2023) for practical reasons (refer to Section 3.4). We use the same testset (i.e., the 8k testset) proposed by LeClair and McMillan (2019). For C/C++, we extract 9k methods for test and 10k methods for training from a dataset proposed by Haque et al. (2021). We used this dataset because it includes 18,775 C/C++ projects and they followed LeClair and McMillan (2019) for preprocessing. For both C/C++ and Java, we used the srcML tool srcml (2025) to generate AST in XML format in keeping with recommendations by Bansal et al. (2023).
Call Graph Generation Call graph is a graph that represents the relationship between each methods (Ryder, 1979). Call graphs is an important information for helping developers comprehend source code (Alanazi et al., 2021; Walunj et al., 2019; Bansal et al., 2024; Alanazi, 2021)Mabakane (2019) also showed that call graph helps to optimize the performance of the code. In this paper, we use Doxygen (doxygen, 2023) to generate call graphs for both C/C++ and Java. Doxygen is a document generation tool that can also generate call graph. For C/C++ dataset, we extract 25k methods as our testset and 20k methods as our training set from (Haque et al., 2021). We generated the call graph within a file for C/C++ to reduce the tokens of the prompt. For Java, we used the testset proposed by LeClair and McMillan (2019) and further extracted 20k methods from the rest of dataset for finetuning. We generated the callgraph from the entire project because Java does not usually have the callgraph within the file. To reduce the tokens of the prompt, we only input the caller methods in the prompt instead of the complete files.
Data flow Graph Generation Data flow graph represents how one variable passes to another method or variable. Data flow graph is important because it helps programmers to do code comprehension (Sihler, 2024; Kargén and Shahmehri, 2012) and increase productivity of software developers (Al-Saiyd, 2017). In this paper, we use Joern (joern, 2025) to generate the data flow of the method. We use Joern because Joern is highly mature for Java and C/C++. For Java dataset, we used the testset recommended by LeClair and McMillan (2019) to avoid data contamination. We further extracted 15k methods other than testset as our training set. For C/C++, we extracted 8k methods as our testset and 8k methods as our training set from (Haque et al., 2021). We use the same procedure to generate data flow for both datasets.
3.4 Code Models
We use two open-source models and two closed-source models in our experiments. Our goal was to use a range of models from a small one where we can control all experimental variables, to large models that represent the state-of-the-art but where we lose control of several variables.
Open Source Models
JAM We used Jam (Su et al., 2023) with 1,024 context length throughout out entire experiments. jam is a language model based on GPT2 architecture for various code intelligence tasks on Java. jam was pretrained with 52m Java methods with clear train/validation/test set, public available dataset, and deduplication toolkit to avoid data contamination. While jam only has 350m parameters, Su and McMillan (2024) showed in their study that it is small enough to be useful for code intelligence tasks. The advantage of this model is that this model can be run on a single consumer hardware, e.g. 16G VRAM GPU and can be used for the experiments with more controls before scaling up to larger language models.
CodeLlaMA Instruct CodeLlaMA instruct (Roziere et al., 2023) is an open-souce model released by Meta. We used CodeLlaMA instruct because it can better understand the prompt. We used the 13B model in our experiments because it can better fit into our 24G GPU (Weyssow et al., 2025) when using Qlora (Dettmers et al., 2024). We used the Qlora released in (Poludin, 2025). CodeLlaMA is the trade-off between closed-sourced models and Jam because we have access to the models although we still do not have the control on the dataset used for pretraining.
Closed Source Models
GPT-4o mini GPT-4o mini is the commercial models released by OpenAI (Achiam et al., 2023). GPT-4o mini is a closed-source model namely we do not have access to the pretraining data and we send our private source code to the model through API. Because of its closed-source, we cannot avoid data contamination issues (Balloccu et al., 2024; Jiang et al., 2024b) and data privacy issues (Wu et al., 2024). Though several studies have shown the success of GPT models (Jain and Purandare, 2025; Wang et al., 2024b).
Gemini Gemini is another commercial model released by Google (Team et al., 2023, 2024) and is enhanced for multimodal capability. Gemini is also a closed-source model. We used Gemini as another closed-souce model because Gemini has been compared with OpenAI’s ChatGPT in several studies in both software engineering (Sobo et al., 2025; Qi et al., 2023; Siam et al., 2024; Su et al., 2024) and other research topics (Qianyi, 2024; Strzalkowski et al., 2024) and has shown superior than GPT in some studies (Lee et al., 2023).
Table 2: Configurations of fine-tuning and evaluation of different models for each RQ. Note, jam is Java-only but for all other models we had separate Java and C fine-tuning and evaluation.
fine-tuning evaluation
model static analysis code dev.
RQ1 jam - code sum. code sum.
- code trans. code trans.
- code gen. code gen.
codellama - code sum. code sum.
- code trans. code trans.
- code gen. code gen.
RQ2 gpt-4o - - ast gen.
- - call graph gen.
- - dataflow gen.
gemini - - ast gen.
- - call graph gen.
- - dataflow gen.
RQ3 jam ast gen. - ast gen.
dataflow gen. - dataflow gen.
codellama ast gen. - ast gen.
call graph gen. - call graph gen.
dataflow gen. - dataflow gen.
RQ4 jam ast gen. code sum. code sum.
dataflow gen. code sum. code sum.
ast gen. code gen. code gen.
dataflow gen. code gen. code gen.
codellama ast gen. code sum. code sum.
call graph gen. code sum. code sum.
dataflow gen. code sum. code sum.
ast gen. code trans. code trans.
call graph gen. code trans. code trans.
dataflow gen. code trans. code trans.
ast gen. code gen. code gen.
call graph gen. code gen. code gen.
dataflow gen. code gen. code gen.
Table 3: Fine-tuning hyperparameters in our experiments.
jam codellama
epoch 3 3
learning rate 3.00E-05 1.00E-04
batch size 4 2
gradient accum. steps 32 16
lora_r - 64
lora_alpha - 16
lora_dropout - 0
quant_type - nf4
bits - 4
3.5 Research Methods – RQ1
Our research method for answering RQ1 is to train each open-source model to do each code development task, then evaluate each model’s performance on that task. Tables 1 and 2 shows these configurations. For example, the first row of Table 2 shows that we train jam to do code summarization, and evaluate it on that same task. The first row of Table 1 show that the dataset we use for code summarization in Java is 170k dataset in Su and McMillan (2024). However, Table 1 also shows that we did not use jam for C/C++ datasets because jam is pretrained for Java only (see Section 3.4). Our guiding principle in all fine-tuning and evaluation is to use the datasets, hyperparameters, and other experimental settings used by papers that introduced those models and datasets; our goal is to understand model capabilities rather than optimizing performance details. We used the base jam and CodeLLaMa models released by their respective papers where those papers pretrained using large datasets of code (see Section 3.4). Then we further trained those models with the training sets which were released with each dataset (see Sections 3.2 and 3.3). We use the hyperparameters recommended in by the papers accompanying each model, which we show in Table 7. Recall from Section 3.4 that we used a LoRA process for CodeLlaMA due to its size, but a full finetune process for jam.
We evaluated each configuration using the metrics and test datasets in the papers for each dataset. For code summarization, we used the metrics METEOR and USE as implemented by Su and McMillan (2024) and recommended by Haque et al. (2022). For code translation and code generation, we used Bleu score (Papineni et al., 2002) as in the dataset that we used for experiments (Iyer et al., 2018; Lu et al., 2021).
3.6 Research Methods – RQ2
Our research method for answering RQ2 is to evaluate closed-source LLMs on static analysis tasks without any finetuning. We compare the results with the tools that we describe in Section 3.3. We show this configuration in Table 2 For example, we show that we evaluate three different static analysis tasks in the first row of RQ2 in Table 2. Our goal here is the same as RQ1. Therefore, we used the default parameters settings throughout the entire dataset although different parameters might have slightly different results.
For C/C++, we used the testset that we extracted to for evaluation. For Java, we used the teset proposed by LeClair and McMillan (2019) for evaluation. For srcML, we used Levenshtein based on (Decker et al., 2020). We used Jaccard similarity as our evaluation metrics for callgraph and data flow graph analysis (Gharibi et al., 2018; Blokhin et al., 2013; Kargén and Shahmehri, 2016). We also computed the pair accuracy and chain accuracy for both tasks.
3.7 Research Methods – RQ3
Our research method for answering RQ3 is to finetune the open-source LLMs and evaluate those models on static analysis tasks. We show those configurations in Tables 1 and 2. For example, we show that we finetune jam to do ast generation and evaluate jam with the same task. We do not finetune jam for callgraph generation because of the context length limit (see Section 3.4). Note that we increase the number of tokens to 10k for callgraph generation because callgraph generation would require larger context size. We use the same hyperparamters and metrics for evaluation as in RQ1 and RQ2.
3.8 Research Methods – RQ4
Our research method for answering RQ4 is to finetune the the open-source LLMs on downstream tasks after those models have been finetuned on static analysis tasks. We show those configurations in Tables 1 and 2. For example, in the first row of jam column in RQ4, we show that we first finetune jam on the AST generation task. Then, we further finetuned it for code summarization. We used the same hyperparameters and metrics described in RQ1.
3.9 Threats to Validity
We organize threats to validity into threats to internal validity, external validity, and construct validity based on Wohlin et al. (2012). The threats to internal validity are that different hyperparamters can have different results. We mitigate this threat by following the accompanying papers. The threats to construct validity are that our experimental results may not be able to generalize to different tasks. We mitigate this by using three different code development tasks and static analysis tasks alone with two different programming languages. The threats to external validity are that different models might generate different results. We mitigate this threat by using two different commercial LLMs and two different open-souce LLMs.
Table 4: Meteor and USE score for code summarization
Metrics
model config METEOR USE p_m p_u
Java jam base 40.95 68.80 - -
srcml 41.28 69.01 0.02 0.06
dataflow graph 40.09 68.32 0.99 0.99
callgraph - - - -
codellama base 51.44 74.56 - -
srcml 51.19 74.48 0.99 0.88
dataflow graph 51.27 74.54 0.94 0.62
callgraph 51.43 74.57 0.53 0.43
C codellama base 29.65 50.62 - -
srcml 29.96 50.85 0.99 0.99
dataflow graph 29.87 50.73 0.99 0.93
callgraph 29.97 50.91 0.99 0.99
Table 5: BLEU score for code generation
model config BLEU p
Java jam base 23.51 -
srcml 24.00 $p<0.01$
dataflow graph 22.79 0.63
callgraph - -
codellama base 31.42 -
srcml 30.40 0.92
dataflow graph 34.22 $p<0.001$
callgraph 28.93 0.99
C codellama base 48.20 -
srcml 48.18 0.70
dataflow graph 47.54 0.50
callgraph 47.82 0.73
Table 6: BLEU score for code translation
model config BLEU p
Java jam base 50.34 -
srcml 49.26 0.90
dataflow graph 50.25 0.08
callgraph - -
codellama base 80.36 -
srcml 80.52 0.56
dataflow graph 79.82 0.57
callgraph 81.84 $p<0.01$
C codellama base 37.82 -
srcml 37.50 0.84
dataflow graph 41.64 0.91
callgraph 39.70 0.91
4 Experimental Results
In this section, we discuss our experimental results to each of our RQs.
4.1 RQ1: Open Models, Code Dev. Tasks
We showed the results for RQ1 in Table 45, and  6. In Java, we found that jam has 40.95 METEOR score and 68.80 USE score for code summarization. Larger models with larger training data further improve the results across three different datasets. For example, we found that CodeLlaMA has 25% improvement in METEOR score and 8% improvement over USE in code summarization. To our surprise, code translation has over 50% improvement when we finetuned it with 13B parameter models. Although the result shows that larger LLMs tend to have the better results compared with smaller models (Zhang et al., 2024a), we cannot attribute the improvement on CodeLlaMA with the size of the model because of the data leakage issues in the training set of those large language models (López et al., 2024). However, we have confidence that those methods are not in the  jam model’s training set.
For code summarization and code translation, we observed the higher performance in Java and low performance in C/C++. The rationale for the higher performance on Java may be because of higher amount of training data as a result of the popularity of the programming language (Jiménez, 2024). Another possible explanation for higher performance on the code translation task in Java is that the syntax of Java and C# are similar to each other and models have more information from source language. However, C and Rust have larger difference in terms of syntax (Shiraishi and Shinagawa, 2024). For example, to define an integer $a=0$ in Java and C#, we can simply write “int a = 0;” in both languages. However, for C and Rust, we would write “int a = 0;” and “let a = 5;”. Overall, these findings showed that LLMs can perform better on tasks with similar syntax and more popular programming languages but perform worse on tasks with different syntax.
In terms of code generation, we observed the low performance in Java and higher performance in C/C++. The possible explanation is that LLMs struggles with code generation when it comes the code with more complicated logic (Zhong and Wang, 2024; Qiu et al., 2024). This result also aligns with recent findings that LLMs usually struggle with generating code on object-oriented programming languages (Jiménez, 2024; Cao et al., 2024). The cause of this difference may be because the inability of LLMs to infer the information from longer context although the context windows is long, e.g. 128k tokens for GPT or Gemini (Li et al., 2024a).
Table 7: Fine-tuning hyperparameters in our experiments.
jam codellama
epoch 3 3
learning rate 3.00E-05 1.00E-04
batch size 4 2
gradient accum. steps 32 16
lora_r - 64
lora_alpha - 16
lora_dropout - 0
quant_type - nf4
bits - 4
Table 8: Results for srcml; gpt-base means the result without in-context learning; gpt-in-context means the result with in-context learning; codellama-finetuned means the result after finetuned.
Metrics
task model Levenshtein
Java srcml gpt-base 0.45
gpt-in-context 0.80
gemini-base 0.64
gemini-in-context 0.93
codellama-finetuned 0.77
jam-finetuned 0.89
C srcml gpt-base 0.46
gpt-in-context 0.60
gemini-base 0.65
gemini-in-context 0.85
codellama-finetuned 0.61
jam-finetuned -
Metrics
task model Jaccard similarity pair accuracy chain accuracy
Java callgraph gpt-base 0.121 0.123 1.10E-04
gpt-in-context 0.122 0.123 6.45E-06
gemini-base 0.186 0.190 1.90E-04
gemini-in-context 0.191 0.195 1.20E-04
codellama-finetuned 0.910 0.940 2.20E-01
dataflow graph gpt-base 0.001 0.004 0
gpt-in-context 0.011 0.029 1.29E-06
gemini-base 0.025 0.007 0
gemini-in-context 0.010 0.036 2.33E-04
codellama-finetuned 0.623 0.771 1.24E-01
jam-finetuned 0.550 0.780 2.50E-01
C callgraph gpt-base 0.160 0.200 8.00E-06
gpt-in-context 0.217 0.280 0
gemini-base 0.180 0.210 6.47E-05
gemini-in-context 0.265 0.312 3.56E-05
codellama-finetuned 0.310 0.550 1.23E-04
dataflow graph gpt-base 0.0009 0.003 1.00E-04
gpt-in-context 0.004 0.010 4.00E-03
gemini-base 0.0016 0.0035 1.35E-03
gemini-in-context 0.007 0.017 6.00E-03
codellama-finetuned 0.120 0.217 8.00E-02
jam-finetuned - - -
Table 9: Results for callgraph and dataflow graph analysis.
(a) Source code
(b) SrcML generated by tool
(c) SrcML generated by GPT without in-context learning
(d) SrcML generated by GPT with in-context learning
Figure 1: Example of SrcML generated by GPT and tool for Java
4.2 RQ2: Closed Models, Static Analysis Tasks
We showed the results for AST generation for both Java and C/C++ in Table 8. In Java, we found that gpt-base has the Levenshtein score of 0.45 and gemini-base has the Levenshtein score of 0.8. We also found the similar results in gemini although we found higher results in gemini, which has Levenshtein score of 0.64 for gemini-base and Levenshtein score of 0.93 for gemini-in-context. For C/C++, we observed the similar trend as in Java that in-context learning helps to improve the overall results on AST generation. However, we observed the smaller improvement in C/C++ dataset. For example, we found around 80% improvement between GPT-base and GPT-in-context in Java dataset. However, we only found around 30% improvement between GPT-base and GPT-in-context in C/C++ dataset, which has around 50% difference in improvement. This result aligns with the finding that LLMs can have better performance on more popular programming languages (Jiménez, 2024). The gap between C/C++ and Java may imply the lack of reasoning skills on LLMs to generalize from one programming language to another programming language.
We showed an example from GPT in Figure 1 for Java. We found that without in-context learning, the model tends to start the AST with <class><name> ClassName</name> even if we only use a single method as we showed in Area 1 in Figure 1(c). With in-context learning, the method can correctly generate the AST that starts with <function> as shown in Area 1 in Figure 1(d). A possible explanation is that LLMs have been trained on huge dataset that has complete class, so those models assume all Java methods are inside a Java class. It is also possible that LLMs just try to learn the correct tags instead of correct syntax as in Area 2 in Figure 1(d) and 1(b).
We showed the results for callgraph and data flow graph in Table 9 for both Java and C/C++. For callgraph generation in Java, we found that gpt-base has Jaccard similarity of 0.121 and gpt-in-context has Jaccard similarity of 0.122. gemini-base and gemini-in-context have slightly higher performance, which is 0.186 and 0.191 respectively. We also found that chain accuracy is the lowest among all other metrics. In C/C++, we found the similar trend that in-context learning does not help to improve the results and chain accuracy is the lowest among all other metrics. For data flow generation, we found the similar results on both Java and C/C++ dataset. Overall, we found that LLMs do a poor jobs on callgraph generation and data flow graph generation and in-context learning does not help to improve the performance for callgraph generation and data flow generation. The possible explanation is that LLMs do not have the fundamental ability for callgraph generation and data flow generation. It is also possible that more samples would help LLMs to learn more useful information for these two tasks (Xu and Zhang, 2024). Another possibility is that simply throwing everything into one giant prompt would not help to improve the model for more complex tasks (Zamfirescu-Pereira et al., 2023) although LLMs can take very long context.
Overall, we found that closed-source LLMs do not do static analysis very well. To our surprise, we found that in-context learning only helps to improve AST generation but not callgraph generation and data flow generation. The possible explanation is that it is easy for LLMs to learn the tags for AST generation. However, LLMs struggle with looking at larger context for callgraph generation or more complex thought process like human programmers would do for data flow analysis. These results suggest that LLMs do not have the same reasoning skill as human programmers.
(a) Source code
(b) SrcML generated by CodeLlaMA
(c) SrcML generated by tool
Figure 2: Example of SrcML generated by CodeLlaMA and tool
4.3 RQ3: Open Models, Static Analysis Tasks
Table 8 showed results for AST generation in Java and C/C++. In Java dataset, we observed that open-source models jam reached Levenshtein score of 0.89 and CodeLlaMA reached Levenshtein score of 0.77 for AST generation. This result shows that LLMs learn meaningful information for AST generation during finetuning. We found that the result of CodeLlaMA is slightly lower. The possible explanation is that we used less data to finetune the model due to the computational cost although CodeLlaMA is already a larger model with more pretrained data compared with jam. To our surprise, we found that jam even outperforms gpt-in-context in this task although we cannot strongly claim that smaller models can do better than gpt-in-context on this task because we did not try to find the best prompt for the task. However, we found that an appropriate finetune helps to improve this task. For C/C++ we also observed the lower Levenshtein score compared with Java. This result aligns with the results from code development tasks that C/C++ has lower performance compared with Java in Section 4.1. Interestingly, we observed it is possible that LLMs make a very simple syntactic error. For example, we observed in Figure 2(b) in Area 1 that the AST generated by CodeLlaMA misses one “)” compared with Area 1 in Figure 2(c). Moreover, we observed in Figure 2(b) in Area 2 that LLMs might continue generating the token at the end. Overall, this shows the lack of structural information in the LLMs.
We showed the result for both C/C++ and Java for callgraph generation in Table 9. We found that CodeLlaMA has high Jaccard similarity score and pair accuracy for callgraph generation after finetuning. However, we found that the chain accuracy is still very low. This result shows that models can only partially generate the correct callgraph. We also found that C/C++ dataset has lower score compared with Java dataset. The possible explanation is that we input the entire file for C/C++ dataset and this makes the problem harder. In addition, we only use 10k tokens during prediction due to hardware limitation. It is also possible that using entire files would require more tokens. However, the overall conclusion that lower chain accuracy and higher pair accuracy and Jaccrad similarity would remain the same based on our findings from Java. Overall, we found that LLMs struggle with looking at the longer context although the input size is long (Li et al., 2024a).
The results for both C/++ and Java for data flow graph generation are in Table 9. Overall, we observed that the results for data flow graph generation has the similar results as in callgraph generation i.e. low chain accuracy but higher Jaccard similarity and pair accuracy and C/C++ dataset has lower results compared with Java. In Java, we also found that jam has very close results with CodeLlaMA on pair accuracy with the same finetuning data on data flow graph generation in Java dataset. This is to our surprise because larger models usually have bigger improvement in code development tasks e.g. 25% improvement on meteor score for code summarization in Java. However, it only shows a slight improvement or even worse on static analysis tasks. This result implies that for static analysis tasks, LLMs tend to saturate faster than other code development tasks. A possible explanation is that it is harder for LLMs to learn the useful information in static analysis tasks.
To sum up, we found the similar results on both open-source models and closed-source models. Specifically, we found that LLMs have larger improvement on AST generation after finetuning. However, LLMs struggle on callgraph generation and data flow graph generation. This result shows LLMs’ inability on more complex reasoning tasks such as static analysis tasks.
Table 10: Code summarization example for Java
Method ID 29591858
static public String makeName( List axes) {
StringBuffer buff = new StringBuffer();
for (int i=0; i<axes.size(); i++) {
CoordinateAxis axis = (CoordinateAxis) axes.get(i);
if (i>0) buff.append("-");
buff.append( axis.getName());
return buff.toString();
codellama-base takes a list of CoordinateAxis objects and returns a string that concatenates their names separated by hyphens
codellama-srcml takes a list of CoordinateAxis objects and concatenates their names with hyphens in between, returning the resulting string
codellama-callgraph takes a list of CoordinateAxis objects and concatenates their names with hyphens in between, returning the resulting string
codellama-dataflow takes a list of CoordinateAxis objects and concatenates their names with hyphens in between, returning the resulting string
reference takes a list of axes and returns a concatenated string of their names joined by ”-”
Table 11: Code summarization example for C/C++
Method ID 120
static void knl_show_mc_route(u32 reg, char *s)
int i;
for (i = 0; i < KNL_MAX_CHANNELS; i++) {
s[i*2] = knl_get_mc_route(i, reg) + ’0’;
s[i*2+1] = ’-’;
s[KNL_MAX_CHANNELS*2 - 1] = ’\0’;
codellama-base show the routing for a given MC channel
codellama-srcml show the multicast route for a given register
codellama-callgraph show the routing for a given MC channel
codellama-dataflow show the routing for a given MC channel
reference Render the MC_ROUTE register in human readable form
Table 12: Code generation example for Java.
Method ID 988
boolean function(ProjCoordinate arg0, double arg1 ){
double loc0 = arg0.x - this.x;
if ( loc0 > arg1 ) { return false; }
else { return true; }
jam-base boolean function( double arg0){ return (( arg0 >= this.x ) && ( arg0 <= this.x + this.z) && ( ( arg0 >= this.y ) && ( arg0 <= this.
jam-srcml boolean function(double arg0, double arg1 ){ return (arg0 >= x) && (arg0 <= arg1); }
jam-dataflow boolean function(Coordinate arg0 ){ return ( x == arg0.x && y == arg0.y && z == arg0.z ); }
Table 13: Code generation example for C/C++.
Method ID 10
int binaryDec(int arr [], int n){
int ans = 0;
for(int i = 0 ; i < n ; i ++ )
ans += arr [ i ] << ( n - i - 1 );
return ans;
codellama-base int toDecimal(int arr[] , int n) { int res = 0 ; for(int i = 0; i < n;i ++){ res += arr [i] * ( 1 << i ); } return res; }
codellama-srcml int toDecimal(int arr [], int n){ int res = 0; for(int i = 0; i < n; i ++ ) res += arr [ i ] * ( 1 << i ); return res; }
codellama-dataflow int toDecimal(int arr [], int n){ int res = 0; for(int i = 0; i < n; i ++){ res += arr [i] * ( 1 << i ); } return res; }
codellama-callgraph int convertFromBinary(int arr []){ int n = 0 , i = 0 ; while(arr[i] != 0 ){ n = n * 2 + arr [i]; i ++; } return n; }
Table 14: Code generation example for C/C++.
Method ID 10
int EnneacontahexagonNum(int n){
return (94 * n * n - 92 * n) / 2;
codellama-base int enneacontahexagonNum(int n){ return ( 118 * n * n - 106 * n ) / 2; }
codellama-srcml int enneacontahexagonNum(int n){ return ( 11 * n * n - 10 * n ) / 2; }
codellama-dataflow int enneacontahexagonNum(int n){ return (11 * n * n - 10 * n) / 2; }
codellama-callgraph int enneacontahexagonNum(int n){ return (14 * n * n - 14 * n + 1) / 2; }
4.4 RQ4: Open Models, Static Analysis + Dev.
We showed the results for code summarization in Table 4 in Java and C/C++. We do not strongly find the statistical difference between the models pretrained with static analysis and the baseline model for both jam and CodeLlaMA for Java and CodeLlaMA for C/C++. The possible explanation is that the model looks at the keyword in the method for source code summarization instead of using static analysis tasks like what human programmers would do to comprehend source code. We showed an example in Java in Table 10. We found that LLMs tend to copy the technical jargon from the method for code summary. Specifically, we found that the summary from CodeLlaMA use “CoordinateAxis” in the summary even after static analysis tasks. Similarly, we also observed the same phenomenon in C/C++. For example, in Table 11, we found that human reference uses “render” as a verb to describe the method. However, the summary from CodeLlaMA uses “show” as a verb, which matches the signature of the function. Interestingly, we found that some of the summaries are the same. This is because we use very small temperature i.e. 1e-9 and models may generate the same result when the weight are similar (Peeperkorn et al., 2024), which further showed that pretraining models with static analysis tasks does not improve code summarization tasks significantly.
Table 5 shows the result for code generation tasks for both Java and C/C++. Overall, we do not strongly observe any statistical difference in Java and C/C++ dataset, i.e. the statistical significant difference across all static analysis tasks and two different programming languages. The possible explanation is that LLMs do not use static analysis tasks as the thought process. For example, we found in Table 4.3 that code generated by jam-srcml has better syntactic structure compared with jam-base. However, the underlying logic is still different from the reference. Another example in C/C++ dataset in Table 4.3 is that we observed that the code generated by CodeLlaMA without pretraining with any static analysis tasks has the logical error in the “toDecimal” method. After pretraining with static analysis tasks, we still found the same logical error. Interestingly, we observed that LLMs might generate a code with incorrect details. In Table 4.3, we found that CodeLlaMA does not generate the correct parameters for Enneacontahexagon numbers. For human being to do this task, they would look at the correct parameters before writing code. This implies that LLMs use a random number or a specific number from other information for some tasks. This shows the lack of reasoning skills in LLMs and LLMs do not use static analysis tasks for code generation.
We showed the results on code translation for both Java and C/C++ in Table 6. In Java, we do not observe statistical difference between the model pretrained with data flow analysis and srcml for both jam and CodeLlaMA although we do observe callgraph helps to improve the result on this task. However, we do not observe any statistical difference between in C/C++ dataset although we do observe the improvement when we pretrained the model with callgraph and data flow graph. The possible explanation is that pretraining with callgraph or data flow helps models look at the larger context and code translation would require a larger context for more accurate translation. However, we cannot strongly attribute the improvement to this because models only have partial improvement on callgraph and data flow generation. Another finding is that code translation have higher performance compared with another task in Java dataset. It is also possible that Java and C# have very similar API calls and syntax, so LLMs can easily copy the information from the source language as we discussed in RQ1 (refer to Section 4.1) as we do not observe similar trend in C to Rust dataset.
Overall, we do not strongly find that static analysis tasks lead to better results on coding tasks for both C/C++ and Java although we do find statistical significance in some cases. These lead to two conclusions 1) LLMs might have different thought process for coding tasks compared with human programmers 2) the high performance might suggest that LLMs use semantics information instead of syntactic information or more complex reasoning.
5 Conclusion
In this paper, we have three key observations. First, for code intelligence tasks, we observed that LLMs usually have higher performance on more popular programming languages. Specifically, we observed models usually have higher performance on Java compared with C/C++. The gap between these two programming languages may be because that Java has more training data (López et al., 2024). Moreover, we also found that the performance of LLMs is usually lower when the underlying structure of source code is more complicated. This shows that LLMs’ reasoning skill is limited.
Second, we found that in-context learning and finetuning only improve the performance on AST generation. However, in-context learning does not improve more complex static analysis tasks and finetuning only improves the more complex static analysis tasks partially. Specifically, we do not observe any significant improvement on callgraph generation and data flow graph generation with in-context learning. Moreover, we found that finetuning improves pair accuracy and Jaccard similarity although we do not observe the significant improvement on chain accuracy. This result shows that LLMs struggle with more complex reasoning.
Finally, we found that pretraining models with various static analysis tasks does not necessarily improve the code intelligence tasks. Specifically, we do not strongly observe the statistical difference across all datasets and all programming languages when we pretraining models with different static analysis tasks. Another evidence that showed that static analysis tasks do not improve code intelligence tasks is that LLMs tend to make the same logical errors after pretraining models with static analysis tasks. To sum up, this shows that LLMs have different thought process compared with human programmers.
Acknowledgements.
This work is supported in part by the NSF grants CCF-2100035, CCF-2211428, and CCF-2211429. Any opinions, findings, and conclusions expressed herein are the authors’ and do not necessarily reflect those of the sponsors.
Ethical approval
This study was conducted in accordance with institutional guidelines and approved by the Institutional Review Board.
Informed consent
Not applicable.
Author Contributions
Chia-Yi Su designed the experiments, implemented the code for the experiments, and wrote the manuscript. Collin McMillan wrote the manuscript and designed the experiments.
Code / Data Availability Statement
We release all data, scripts, survey materials, and results at https://github.com/apcl-research/llm-reason.
Clinical trial number
Not applicable
Conflict of interest
The authors declare that they have no conflict of interest.